Exercises with the nuclear shell model, spin and parity .(8) Define the azimuthal angle e associated to each azimuthal quantum number m in the cases below (j-1/2, 1, 2) can you show me step by step how to get to these answers 1/2 n/2 2n #54.7,-54.7 deg 45., 90, -45 deg 35.3, 65.9, 90,-65.9,-35.3 deg
Answer #1

For every value of  j , we will have 2j+1 different value of m_j, ranging from -jldots,0,ldots,+j

Total angular momentum will be j(j+1)

The Z-component of total angular momentum will be mi h

The different azimuthal angles are given by 7m θ = cos = cos

i) For j=rac{1}{2} ,  2 mi,   154.70 θ cos COS

For j=rac{1}{2} ,  2 mi- , azimuthal angle is θー-54.70

ii)

For j=1  ,  mi ,   θ os-1 -1 = 45° = COS 2

For j=0  ,  m_j=0 , heta=cos^{-1}left ( rac{0 }{sqrt{1(1+1)}} ight )=90degree

For j=1  ,  m_j=-1 ,  θ 45° ー _

iii)

For j=2  ,  m_j=2 ,   heta=cos^{-1}left ( rac{2 }{sqrt{2(2+1)}} ight )=cos^{-1}left ( rac{2}{sqrt{6}} ight )=35.3degree

For j=2  ,  mi ,   heta=cos^{-1}left ( rac{1 }{sqrt{2(2+1)}} ight )=cos^{-1}left ( rac{1}{sqrt{6}} ight )=65.9degree

For j=2  ,  m_j=0 ,   θ cos-1 COS -90°

For j=2  ,  m_j=-1 ,   θ -65.9.

For j=2  ,  m_j=-2 ,   heta=-35.3degree​​​​​​​